# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

### v4

@@
Abstract: The embrace of globalization and protectionism among economies has ebbed and flowed over the past few decades. These fluctuations call for quantitative analytics to help countries improve their trade policies. Changing attitudes toward globalization also imply that the best trade policies vary over time and are country-specific. We argue that the imports and exports of all economies constitute a counterbalanced network where conflict and cooperation are two sides of the same coin. Quantitative competitiveness is then formulated for each country using a network counterbalance equilibrium. A country could improve its relative strength in the network by embracing globalization, protectionism, trade collaboration, or conflict. This paper presents some necessary conditions for globalization and trade wars, evaluates their side effects, derives national bargaining powers, identifies appropriate targets for conflict or collaboration, and recommends fair resolutions for trade conflicts. Data and events from 2000 to 2019 support these conditions.
Keywords:  globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma
JEL Code: C71, C78, F11, F13, F15, O24
@@ 1 Introduction
De-globalization and a trade war are two levels of economic conflict; the latter directly involves only two trade partners, whereas the former represents one country’s disintegration from the global economy. Nevertheless, both could adversely affect global economic activity. This paper applies a unified game-theoretic approach to formulate simple but decisive rules to mitigate either type of conflict. We focus on import and export activities, ignoring political, territorial, ideological, military, and cultural conflicts. However, geopolitical data have already been incorporated into trade behaviors and volumes. From this perspective, a country’s primary purpose in engaging in trade wars or globalization is to improve its competitiveness and trade balance in the trade system. Also, the engagement affects the future geopolitical landscape.
@@
To analytically examine trade wars and globalization, we face several daunting challenges. First, there are always conflicts and cooperation in international trade. While the trade of goods and services results from bilateral collaboration, a country’s growing strength could weaken its trade partner’s position. Thus, a purely cooperative or non-cooperative study could be too impotent to capture the complexity of trade relationships. The second challenge is adequately constructing a high-level description of the interactive trade system because exports and imports have already reflected the comparative advantages of specific goods and services. The system is a directional network from which we aim to quantitatively formulate the economies’ utility or social welfare functions in this conflicting and cooperative context. Next, policy variables, such as imports and exports as production’s percentages, should be linked to these economic objectives. So, countries could apply optimization methods to choose appropriate policies to achieve their economic goals. Finally, it is fruitful to bring data into the analytics to explain past events and find optimal solutions for the near future. The solutions could identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization. In the literature, studies on globalization often capitalize on econometric analysis, which uses historical data to estimate unknown parameters. However, the estimated effects could hardly catch up with the recent changes or seize both cooperation and non-cooperation.
We measure the relative strengths of the countries in the network by the solution to the counterbalance equilibrium, defined by Hu and Shapley, (2003), and use it as an objective in our analysis. The solution, called authority distribution, integrates both cooperation and competition inside the global economic system. The data-driven distribution is a social welfare function and an endogenously weighted average of utility functions. Moreover, it measures the competitive advantages of the countries and their long-term influences on the whole system. Also, this competitiveness brings extra welfare beyond the economy of scale and comparative advantage. The additional advantage, called the Matthew effect, is another driving force for competition after comparative advantages.
This paper attempts to provide strategic solutions for policymakers and ignores how an individual firm or economic agent improves firm’s competitiveness through research and development. Maximizing national competitive strength is a purview of government policymakers rather than trade companies, which often capitalize on comparative advantages. A country is not merely an aggregate of firms, and the interests of trade companies only sometimes align with those of policymakers, who also consider the side effects of trade, such as the shrinkage of domestic labor markets and the overseas shift of industries. So, our studies neglect the firm-based micro foundation and maximize a social welfare function, pivoting domestic production decision-making. However, the competitiveness measure is also a weighted average of comparative advantages that considerably reflect the profits of domestic firms engaging in international trade and the welfare of consumers.
When starting a trade war against a trade partner, a country raises tariffs or creates other barriers to increase competitiveness. However, the result depends on the counterpart’s reaction. We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in a conflict or cooperation. On the other hand, the reaction may also benchmark their production ratio to avoid trade deficits. Using the threshold and the benchmark ratios, we derive a Nash bargaining solution for countries with heterogeneous preferences between competitiveness and trade balance, and the solution generates consistent national bargaining power. It provides a solution to resolve bilateral trade frictions.
A country with no specific targets to confront could choose between further globalization or protectionism. By minor perturbations in its overall exports, we calibrate the effect on its competitiveness from the counterbalance equilibrium. The sign of the effect tells if the country should go for more globalization or not. There are many debates over economic globalization in the literature, and Stiglitz (2002, 2017) documented their detailed viewpoints. Different viewpoints, assumptions, and considerations could reach opposite conclusions about globalization for a country. Our data-driven approach ignores arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security. Trade policymakers may carefully weigh these considerations, and the trade data may already reflect them. Hence, a country could engage in further globalization or protectionism, depending on its particular situation at a specific time. By analyzing the ComTrade data of the United Nations (2021) for the years 2000 through 2019, we find that protectionism could have been a better choice for the USA and the United Kingdom but a worse option for China, Germany, Japan, and Russia if they targeted competitiveness and trade balances only. Data from 2017 also show that trade wars could be an option for the USA to maintain its competitiveness. However, adverse side effects on third parties and allies were also tangible. According to our theorem of impossibility trilemma, it has to reduce imports if it aims to lower its national debt by earning trade surpluses and also preserve its competitiveness.
In the following exposition, we first introduce the competitiveness objective in Section 2. Then, Section 3 formulates a necessary condition for a country to launch a trade war or form a trade partnership to optimize its objectives. After that, a bargaining solution is derived in Section 4 to resolve conflicts or share profits. Next, Section 5 discusses a similar condition for globalization or protectionism. Then, in an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019. Finally, we discuss several extensions and limitations of this framework in Section 7. Our exposition is self-contained and the proofs are in the Appendix.
2 Competitive Advantage in the International Trade System
Based on Adam Smith’s economy of scale and David Ricardo’s comparative advantage, if all production factors were free of flow, then, hopefully, global production would be maximized. However, not all countries share the same well-off status; some succeed over some time, whereas others fail during the same period. Even the most powerful economies have failed in human history, and there needs to be more mechanisms to enforce the fairness of profit sharing in international trade. Thus, a country has to compete for its share and, at the same time, cooperate with others to maximize global production. Both competition and cooperation are essential in international trade.
Domestic production is paramount for policymakers, whereas profit maximization is for firms or producers. Production with zero or slightly negative profit means nothing to firm owners but may still be an option for the government to uplift employment. Production maximization, for example, results in less unemployment, while labor may be cut or resourced overseas to maximize shareholders’ profits. Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors to the firms. Furthermore, the government not only regulates profit-maximization behavior but also improves consumer welfare and redistributes the profits of domestic firms through taxation.
@@ 2 Competitive Advantage in the International Trade System
We regard the international trade system as a game on a network in which countries act as nodes and trades as directional edges. The trade volumes flow along the edges, and imports and exports move in opposite directions. In the multilateral network, a unilateral action could achieve unpredictable or even opposite results. Trade wars or antiglobalization are often framed to shield domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping. These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005). Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of the trade counterparts.
@@
This section introduces a new type of competitiveness as an objective and social welfare function for the countries in the network. Competitiveness is a relative measure, and we use a constant-sum game to derive it. Absolute production and total trade volume, thus, may give limited information on the economy’s competitive position. The game mocks the in-and-out flows over the network; without data change, these steady flows reach a static equilibrium due to long-run accumulation. The equilibrium is the competitiveness vector, which describes the utilities of the countries’ trade activities. We detail the steps toward the data-based utility and social welfare functions, which are unfamiliar in the trade literature.
2.1 Long-Run Influence in International Trade System
@@
For countries in the network, we describe their aggregate trade activities with a square matrix $P$. Assume there are $n$ countries labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$. For any $i,j\in\mathcal{N}$ with $i\not=j$, let $P_{ij}$ be the fraction of country $i$’s production, which exports to country $j$. The exports include all services and goods, such as final goods and intermediate goods, both of which may also contain parts made by third parties or country $j$. Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as domestically produced final products. Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are attributable directly to the production, such as raw materials, intermediate goods, and crude oil. It also includes exports of non-final products but excludes relevant imports, which are already included in $g_{i}$. Besides, $g_{i}$ should exclude exported final goods because they are already included in the GDP. Also, imported final goods which are not directly used in the domestic production should not be counted in $g_{i}$. We use $P_{ii}=1-\sum\limits_{j\not=i}P_{ij}\geq 0$ for the non-exporting fraction. Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $j\in\mathcal{N}$. We place all the fractions $P_{ij}$ into a $n\times n$ stochastic matrix $P=[P_{ij}]$.
@@
We look for a utility function which determines country $i$’s production or expected production from a demand- or consumption-side perspective. For any $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
@@
together with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$. So, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$. It is normalized since $v_{i}(\mathcal{N})=1$. Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game and its Shapley value is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$ (Shapley, 1953). Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern (1953) utility function, i.e., $P_{ij}$ is $j$’s utility in the game of $(\mathcal{N},v_{i})$. So, countries have different preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$. However, their preferences change as $i$ varies in $\mathcal{N}$. In the matrix $P$, therefore, each row is an individual utility function whereas column countries are the arguments or social states of the function. A social welfare function generally combines these individual utility functions into one, accounting for the diverse judgments of every country in $\mathcal{N}$. It rates the set of countries by $\mathcal{N}$ as a whole.
The element $P_{ij}$ is a type of pro rata power when country $i$ makes its production. The multilinear extension of $v_{i}$ is defined by
@@
where $x_{j}\in[0,1]$ is the probability with which $j$ participates in the consumption-driven production. Clearly, when restricted to $\{0,1\}^{n}$ or the corners of $[0,1]^{n}$, $\tilde{v}_{i}$ reduces to $v_{i}$. When the countries act independently, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ is country $i$’s expected production. Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also means $j$’s progress in completing its imports from $i$. In addition, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$. Without any prior knowledge about the participation probabilities or progresses, we assume they advance evenly along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$. For an annual production, for example, each country finishes $1/12=8.33\%$ of its imports from $i$ in any month. Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s influence or power when country $i$ completes its production from 00 to $1$. According to Owen, (1972), the integral equals the Shapley value for the coalitional game, i.e.,
@@
where $\top$” denotes vector or matrix transpose.
From a supply-side perspective, inputs to the production function in Eq. (1) can include any variables that affect the production and services of country $i$. However, it is impossible to compile a complete list of these variables. Their effects are also time-varying and country-specific; each product has a unique way of combining them together. For example, inflows of financial assets, including remittances and foreign direct investments, are influential in making domestic products. They are primarily payments for past, current, or future exports. Trade is about exchanging goods and services; as the medium of exchange, monetary flow makes transactions possible in modern international trade. Nevertheless, payments alone could not fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only. Zero payments may occur for intrafirm transactions of exports within a multinational company. Also, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively. Besides, a country can accumulate trade deficits if it is a highly desirable destination for foreign investment. For example, U.S. Treasury Bonds are lucrative and low-risky investment instruments, so foreign governments or companies must earn extra U.S. dollars from their trades to put into their foreign investment portfolios.
The direct influences in matrix $P$ automatically generate indirect ones in the closed global trade system. In a chain of production, even though the final product bears the mark of being made in one country, its components or parts may come from elsewhere. When China exports a smartphone to the USA, for a fictitious example, the chip and design software of the phone could come from the USA and the camera from Korea. Japan could manufacture the camera screen using US patents. To integrate these spillovers, we consider indirect and long-term impacts by the powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$. In general, $P^{t}$ contains the $t$-step aggregate impacts of all supply chains that crisscross the globalized world. For example, its value at the $i$th row and $j$th column, $\left(P^{t}\right)_{ij}$, accesses all the countries by $t-1$ times because
@@
in which $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$. It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs all countries in $\mathcal{N}$.
Under some general conditions (i.e., aperiodicity and irreducibility), the rows of $P^{\infty}$ are a constant row vector $\pi$, which exists uniquely. These conditions are obviously satisfied in the closed and well-connected trade system. Since $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ regardless of the choice of $i$, $\pi_{j}$ measures country $j$’s long-run influence on the whole trade system. The direct influences in $P$ pass through interlinked global value chains and there are millions of them. As a result, $P^{t}$ converges only if $t$ is large enough for all the direct impacts to travel from the top of global economic chains down to the very bottom. The existence and uniqueness of $\pi$ can be found in the theory of Markov chains (e.g., Karlin and Taylor, 2012). The limit of $P^{t}$ provides a computational method to find $\pi$, and its speed of convergence is linear.

### v5

@@
Abstract
This paper explores the exciting dynamics of international trade, focusing on the strategic balance between competition and cooperation in the global trade network. It highlights how competitive advantages, rather than comparative advantages, drive trade conflicts and deglobalization. By using the balance of power, the paper introduces a quantitative measure of competitiveness that serves as a common goal for all countries, alongside trade balance. It then examines how nations can boost their competitive strength and trade balance through globalization, protectionism, trade collaboration, or conflict. The study provides practical insights for policymakers on managing trade relationships, resolving conflicts, and determining the right level of globalization by analyzing trade data. The analysis is supported by a comparison between theory-driven quantitative evidence and historical events, using real-world trade data from 2000 to 2019.
keywords:
globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma JEL Code: C71 , C78 , F11 , F13 , F15 , O24
Declarations of interests: I have nothing to declare.
Funding: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Declaration of Generative AI and AI-assisted technologies in the writing process: During the preparation of this work, the author never used any Generative AI and AI-assisted technologies.
@@ 1 Introduction
Adam Smith’s concept of economies of scale and David Ricardo’s theory of comparative advantage suggest that global production would be maximized if all production factors could move freely. However, not all countries are equally prosperous; some succeed while others struggle. Even powerful economies have failed in the past, highlighting the need for fair profit sharing in international trade. Therefore, countries must compete for their share while cooperating to maximize global production. Both competition and cooperation are essential in international trade. This paper offers fair and strategic solutions for each nation within the trade network.
@@
Analyzing trade wars and globalization presents several challenges. First, international trade involves both conflict and cooperation. While voluntary trade of goods and services results from bilateral collaboration, a country’s growing strength can weaken its trade partner’s position. Thus, a purely cooperative or non-cooperative study may not capture the complexity of trade relationships. The second challenge is creating a high-level description of the interactive trade system that focuses on competitive advantage, as exports and imports already reflect the comparative advantages of specific goods and services. The system is a directional network from which we aim to quantitatively formulate the economies’ utility and social welfare functions in this conflicting and cooperative context. Policy variables, such as imports and exports as percentages of production, should be linked to these economic objectives. Countries can then apply optimization methods to choose appropriate policies to achieve their economic goals. Finally, incorporating data into the analytics can explain past events and find optimal solutions for the near future. The solutions can identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization.
@@
We view the international trade system as a game on a network where countries act as nodes and trades as directional edges (e.g., Chaney, 2014; $\ddot{\mathrm{O}}$nder and Yilmazkuday, 2016). Trade volumes flow along the edges, with imports and exports moving in opposite directions. In the multilateral network, a unilateral action could achieve unpredictable or even opposite results. Trade wars or anti-globalization efforts are often framed to protect domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping measures. These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005). Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of trade counterparts. For each country, the strengths and weaknesses lie in its domestic products and foreign imports and exports, forming a coalitional game. The Shapley value (1953) of the game provides a revealed cardinal preference or utility function for the individual country (Roth, 1977) and a power index in completing domestic production.
The common battlefield for competition and cooperation is a social welfare function, which weights the individual utility functions (e.g., Harsanyi, 1955; Hammond, 1987). These weights are chosen to balance the powers within the network (Hu and Shapley, 2003). The social welfare function measures the relative strengths of each country and serves as an objective in our analysis. It integrates both cooperation and competition within the global economic system, assessing the competitive advantages of countries and their long-term impacts on the entire system. This competitiveness generates additional welfare beyond economies of scale. The extra advantage, known as the Matthew effect, is another driving force for collaboration beyond comparative advantages.
De-globalization and trade wars represent two levels of economic conflict. The former involves a country’s withdrawal from the global economy, while the latter involves direct conflict between two trade partners. Both can negatively affect global economic activity. We apply a unified game-theoretic approach to formulate simple yet decisive rules to mitigate either type of conflict. Given that competitiveness has a constant sum in the game, the pursuit of competitiveness becomes the primary reason for initiating a trade war or anti-globalization action, while comparative advantages promote globalization and trade collaboration.
This study focuses on import and export activities, considering that geopolitical factors may already be incorporated into trade behaviors and volumes. Given the rationality of importers and exporters, trade data largely reflect political, territorial, ideological, cultural, war, national security, and other geopolitical considerations, alongside transaction and shipping costs, resource endowment, industrial distribution, and location advantages. From this perspective, a country’s primary objective in engaging in trade wars or de-globalization is to enhance its trade balance and competitiveness, as indicated by the data-driven social welfare function.
The objective of this paper is to offer strategic solutions for policymakers to enhance national competitive strength, rather than focusing solely on the competitiveness of individual firms. Government policymakers, rather than trade companies, are responsible for improving national competitiveness, as trade companies often exploit comparative advantages. A nation is not merely a collection of firms, and the interests of trade companies do not always align with those of policymakers, who must also consider the negative impacts of trade, such as the contraction of domestic labor markets and the relocation of industries overseas. Therefore, our study disregards production-based foundations (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014) and instead maximizes the social welfare function, emphasizing domestic production decision-making by global consumers. However, the measure of competitiveness is also a weighted average of comparative advantages, weighted by the preferences of consumers, whether affluent or impoverished.
When a country initiates a trade war against a trade partner, it raises tariffs or imposes other barriers to enhance competitiveness. However, the outcome depends on the partner’s response. We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in conflict or cooperation. Additionally, the reaction may benchmark their production ratio to avoid trade deficits. Utilizing the threshold and benchmark ratios, we derive Nash bargaining solutions (1950), generating consistent national bargaining power and providing methods to resolve bilateral trade frictions. A country with no specific targets to confront could choose between further globalization or protectionism. By making minor adjustments to its overall exports, we calibrate the effect on its competitiveness. The sign of the effect indicates whether the country should pursue further globalization. The analyses of trade wars and globalization also shed light on forming multilateral trade agreements.
The solutions are data-driven, relying solely on trade and production data. For a specific context with a set of fair assumptions, many researchers have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005). There are also numerous debates over economic globalization in the literature, with various detailed viewpoints documented in Stiglitz (2002, 2017). Different perspectives, assumptions, and considerations can lead to opposing conclusions about globalization for a country. Our data-driven approach disregards arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security. Trade policymakers may carefully weigh these considerations, and the trade data may already reflect them. Hence, a country could engage in further globalization or protectionism, depending on its particular situation at a specific time.
We test our new theories by analyzing ComTrade data from the United Nations (2021) for the years 2000 through 2019, comparing theory-derived results with major economic events during the same period. Our findings suggest that protectionism could have been a more advantageous strategy for the USA and the United Kingdom (UK), but a less favorable option for China, Germany, Japan, and Russia, particularly if their objectives were centered on competitiveness. Data from 2017 indicate that trade wars might be a viable strategy for the USA to maintain its competitiveness and reduce its trade deficits, and we assess the impacts of the China–USA trade war in 2018 and 2019. However, we also observe significant adverse side effects on third parties. According to our theorem of impossibility trilemma, the USA must reduce imports if it aims to lower its national debt by generating trade surpluses while preserving its competitiveness.
Our contribution to the trade literature lies in introducing a novel strategic and comprehensive framework for analyzing cooperation and competition within the global economic system. Our analyses eschew unnecessary assumptions and excessive mathematization, which can lead to misleading conclusions and policy implications (e.g., Romer, 2015). Unlike traditional studies, our approach does not rely on unknown parameters and formulates a quantitative measure of competitiveness for each country using the balance of power. This measure encompasses overall strength, including both direct and indirect influences. The study also derives national bargaining powers and offers strategic solutions for policymakers to manage trade relationships effectively.
In the following exposition, we first introduce the competitiveness objective in Section 2. Section 3 formulates a necessary condition for a country to initiate a trade war or form a trade partnership to optimize its objectives. Section 4 derives Nash bargaining solutions to resolve bilateral conflicts or share profits. Section 5 discusses a similar condition for globalization or protectionism. In an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019. Finally, Section 7 discusses several extensions and limitations of this framework. Our exposition is self-contained, and the proofs are in the Appendix.
2 A Network-based Social Welfare Function
@@
This section introduces a novel type of competitiveness and social welfare function for countries within the trade network. Competitiveness is a relative measure derived from a constant-sum game. Absolute production and total trade volume may provide limited information regarding an economy’s competitive position. The game simulates the inflows and outflows within the network; without change, these steady flows reach a static equilibrium due to long-term accumulation. The equilibrium represents the competitiveness vector, which describes the utilities of the countries’ positions in the network game. We detail the steps toward the data-based social welfare function and its unique properties, which are unfamiliar in the trade literature.
Domestic production is paramount for policymakers, whereas profit maximization is crucial for firms or producers. Production with zero or slightly negative profit may be insignificant to firm owners but can still be an option for the government to boost employment. For example, production maximization results in less unemployment, while labor may be resourced overseas to maximize shareholders’ profits. Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors and the distribution of profit to the firms. Furthermore, the government not only regulates profit-maximization behavior but also improves consumer welfare and redistributes the profits of domestic firms through taxation.
@@
For countries in the network, we describe their aggregate trade activities with a square matrix $P$. Assume there are $n$ countries, labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$. For any $i,j\in\mathcal{N}$ with $i\neq j$, let $P_{ij}$ be the fraction of country $i$’s production that exports to country $j$. The exports include all services and goods, including intermediate goods; they may contain intermediate goods made by third parties or country $j$. Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as all domestically produced final products. Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are directly attributable to the production, such as raw materials, intermediate goods, and crude oil. It also includes exports of non-final products but excludes relevant imports, which are already included in $g_{i}$. Besides, $g_{i}$ should exclude exported final goods because they are already included in the GDP. Also, imported final goods which are not directly used in the domestic production should not be counted in $g_{i}$. We use $P_{ii}=1-\sum\limits_{j\neq i}P_{ij}\geq 0$ for the non-exporting fraction. Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $i\in\mathcal{N}$. We place all the fractions $P_{ij}$ into an $n\times n$ stochastic matrix $P=[P_{ij}]$.
@@
Each row of $P$ defines an individual utility function that determines a specific country’s production from a demand- or consumption-side perspective. For any $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
@@
with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$. Thus, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$. Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game, and its Shapley value (1953) is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$. Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern utility function (1953), i.e., $P_{ij}$ measures an appraisal agent’s satisfaction when they play country $j$’s role in the game of $(\mathcal{N},v_{i})$. The agent is neutral with respect to ordinary and strategic risks detailed in Roth (1977). Thus, countries have cardinal preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$. Yet, their preferences change as $i$ varies in $\mathcal{N}$. In the matrix $P$, therefore, each row is an individual utility function, whereas column countries are the arguments or social states of the function. A social welfare function generally combines these individual utility functions into one, accounting for the diverse value judgments of countries in $\mathcal{N}$. It rates the set of countries in $\mathcal{N}$ as a whole and serves as a common platform for the countries to compete and collaborate.
2.1 Long-Run Influence in the Trade System
With the normalization (i.e., $v_{i}(\mathcal{N})=1$ and $v_{i}(\emptyset)=0$) and monotonicity of $v_{i}$, $P_{ij}$ bears much similarity to the Shapley-Shubik power index (1954), measuring country $j$’s chance of pivoting in forming the result $v_{i}(\mathcal{N})$. Indeed, $P_{ij}$ is a type of pro rata power when country $i$ makes its full production. The multilinear extension of $v_{i}$ is a function defined by:
@@ 2.1 Long-Run Influence in the Trade System
where $x_{j}\in[0,1]$ is the probability with which country $j$ participates in the consumption-driven production. Clearly, when restricted to $\{0,1\}^{n}$ or the corners of $[0,1]^{n}$, $\tilde{v}_{i}$ reduces to $v_{i}$. When the countries act independently, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ is country $i$’s expected production. Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also represents $j$’s progress in completing its imports from $i$. Additionally, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$. Without any prior knowledge about the participation probabilities or progresses, we assume they advance evenly along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$. For example, if country $j$ completes 25% of its consumption, all other countries also finish their 25% at the same time. Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s overall influence or power that propels country $i$ to complete the production. According to Owen (1972), the integral equals the Shapley value for the coalitional game, i.e.,
@@
where $\top$” denotes vector or matrix transpose. Therefore, $P$ contains $n\times n$ quantitative powers, and the balance of power in $P$ refers to a state of stability.
From a supply-side perspective, inputs to the production function in Eq. (1) can include any variables that affect the production and services of country $i$. However, it is impossible to compile a complete list of these variables. Their effects are also time-varying and country-specific; each product has a unique way of combining them. For example, inflows of financial assets, including remittances and foreign direct investments, are influential in making domestic products. These assets are primarily payments for past, current, or future exports. Trade involves exchanging goods and services; and monetary flow as the medium of exchange makes transactions possible in modern international trade. Nevertheless, payments alone cannot fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only. Zero payments may occur for intrafirm transactions of exports within a multinational company. Additionally, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively. A country can accumulate trade deficits if it is a highly desirable destination for foreign investment. For example, U.S. Treasury Bonds are lucrative and risk-free investment instruments after the Cold War, so foreign governments or companies must earn extra U.S. dollars from their trades to add to their foreign investment portfolios.
The direct influences in matrix $P$ automatically generate indirect ones in the closed global trade system. In a chain of production, even though the final product bears the mark of being made in one country, its components or parts may come from elsewhere. When China exports a smartphone to the USA, for example, the chip and design software of the phone could come from the USA, and the camera from Korea. Japan could manufacture the camera screen using US patents. To integrate these spillovers, we consider indirect and long-term impacts by the powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$. In general, $P^{t}$ contains the $t$-step aggregate impacts of all supply chains that crisscross the globalized world. For example, its value at the $i$th row and $j$th column, $\left(P^{t}\right)_{ij}$, accesses all the countries by $t-1$ times because
@@
where $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$. It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs through all countries in $\mathcal{N}$.
Under some general conditions (i.e., aperiodicity and irreducibility), the rows of $P^{\infty}$ are a constant row vector $\pi$, which exists uniquely. These conditions are obviously satisfied in the closed and well-connected trade system. Since $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ regardless of the choice of $i$, $\pi_{j}$ measures country $j$’s long-run influence on the whole trade system. The direct influences in $P$ pass through interlinked global value chains and there are millions of them. Consequently, $P^{t}$ converges only if $t$ is large enough. The existence and uniqueness of $\pi$ can be found in the theory of Markov chains (e.g., Karlin and Taylor, 2012). The limit of $P^{t}$ provides a computational method to find $\pi$, and its speed of convergence is linear.